The following is an incomplete simple frequency distribution table for the number of mistakes made during a series of military combat readiness training exercises.

Number of
Mistakes
Frequency
10–A 1
B–9 3
0–4 C
N = 15

Find the missing values for A, B, and C.

A =
B =
C =
Answer #1

We can use here definition of class width

Class width is difference between lower(or upper) limits of consecutive classes.

9-4=5

Therefore, 10-A = 10-14

and B-9 = 5-9

Total frequency = N=15

Therefore, 15 = 1 +3 + C

=> C=11

A=14

B=5

C=11